1. Chapter 5, Combined Model (4 points): Assume that real output is fixed at 500, consumption is fixed at С-350. government spending is fixed at G-60. ·Assume that the Investment function takes the following form 1(r) = 95-r, where r is given in percent terms (i.e., r-2 signifies an interest rate of 2%) » Assume that the quantity equation of money holds M-V- P.Y and the velocity of money is constant V (a) If the money supply increases by 2%, by what percentage rate do prices change? That is, what is the inflation rate? (b) What is the real rate of investment r in this economy? What is the invest- ment level 1? (c) What is the nominal interest rate i? (d) If government spending G increases to 65, what happens to the real rate of investment r, investment I, and the nominal interest rate i?

Macroeconomics - Inflation rate
Answer #1

a) Here any increase in money supply will have effect on price level as there is fall in value of money due to increase in money supply

Velocity of money is constant = V

Y = 500

Price level = p

New price level = P

Money supply = M

New money supply = M (1+2%) = 1.02M

MV = pY ---------eq2

So new equation

1.02M = PY------eq2

Eq2/eq1 will give

1.02 = P/p

P = 1.02p (price level change)

Inflation rate =( (P/p)-1)*100

Inflation rate = (1.02-1)*100 = 2%

b) Y = C+I+G

500 = 350+95-r+60

500=505-r

r = 5%

I = 95-r = 95-5 = 90

C) Nominal interest rate = real interest rate + inflation rate

Nominal interest rate = 5%+2% = 7%

d) G = 65

C = 350

I = 95-r

Y = 500 = C+I+G

500= 350+95-r+65

r = 510-500 = 10%

I = 95-r = 95-10 = 85

Nominal interest rate = r + inflation rate

Nominal interest rate =10%+2%= 12%